40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 6 01:35:03 2012, kiwamu, Update, LSC, MICH characterization 
    Reply  Tue Mar 6 23:14:38 2012, rana, Update, SUS, Optical Lever had bad filters Untitled.png
Message ID: 6362     Entry time: Tue Mar 6 01:35:03 2012     Reply to this: 6367
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: MICH characterization 

[Keiko / Kiwamu]

 Update on the MICH characterization:

  • The OSAs weren't so great because the 11 MHz sidebands were covered by the carrier's tail
    • It seemed that the frequency resolution depended on the mode matching. We will try improving the mode matching tomorrow.
  • The noise budget looked very bad
    • There were huge peaks at 1 Hz, 3 Hz, 16.5 Hz and 23 Hz. Something is crazy in the vertex suspensions.
    • Keiko will post the calibrated noise budget.
  • The MICH response at AS55Q was measured and we will calibrate it into watts / meter.

 

ELOG V3.1.3-